OpenStudy (anonymous):

What is the slope of a line perpendicular to the line with equation y = 6x + 2?

OpenStudy (anonymous):

A.6 B.-6 C.3 D.- 1/6

OpenStudy (anonymous):

If you know the slope of the first line, then what do you automatically know about the slope of a line perpendicular to it? Use that property to compute the slope of the second line.

OpenStudy (anonymous):

-6x + y = 4 you can move the variables around to get the form: y = mx + b, which is the slope intercept form. y = 6x + 4 here: m = 6, b = 4 m is the slope of the line, so slope is 6 any line parallel to this line will also have the slope of 6 any line perpendicular to the given line will have the slope of: -1/m1 we know m1 = 6, so the perpendicular line will have the slope of: -1/6

OpenStudy (anonymous):

1 sec

OpenStudy (reni_love_ela):

There you go

OpenStudy (reni_love_ela):

So B is your ansswer

OpenStudy (campbell_st):

the answer is not B

OpenStudy (campbell_st):

if you have a line y = mx + b the slope of the line is m the perpendicular line has a slope \[m_{perpedicular} = \frac{-1}{m}\] the product of the 2 slopes \[m \times m_{perpendicular} = -1\] so find the slope of your line.. and then find the negative reciprocal

OpenStudy (anonymous):

thx